# General Chen-Ricci inequalities for Riemannian submersions and Riemannian maps

### v1

@@ 3.3 Applications
For generalized complex space forms
Here, we obtain the Chen-Ricci inequalities for Riemannian submersions whose source spaces are generalized complex space forms $\left({M_{1}}^{m_{1}}\left(c_{1},c_{2}\right),g_{1},J\right)$ using the general inequalities obtained in Subsection 3.2. We also provide these inequalities when Riemannian submersions are geometrically structured and validate them with the existing cases where the source spaces are complex space forms. In this case, for any ${\cal Y}\in\Gamma(TM_{1})$, we write
$J{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ such that $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
@@
Let $F:\left(M_{1}^{m_{1}}\left(c_{1},c_{2}\right),g_{1},J\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized complex space form onto a Riemannian manifold. Then we have
@@
[…]
Proof.
Let $M_{1}$ be a generalized complex space form. Then, using (2.1), we derive
and
Adding equations (3.25), (3.26), and (3.27), we obtain
Then by using (3.25) and (3.11), (3.26) and (3.15), (3.3) and (3.3), we obtain the desired inequalities (3.22), (3.23) and (3.4) respectively. ∎
Remark 3.5.
[…]
From [52, 53, 54, 55, 47, 59], we have the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost Hermitian manifolds onto Riemannian manifolds. These notions impose different-different conditions on the distribution $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the complex structure $J$. Using these conditions directly, we can prove the following result along the lines of the proof of Theorem 3.4.
Theorem 3.6.
Let $F:\left(M_{1}^{m_{1}}\left(c\right),g_{1},J\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a complex space form onto a Riemannian manifold.
  1. $1.$
    If $F$ is an invariant Riemannian submersion (i.e. $J(\ker F_{\ast})=\ker F_{\ast}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)+\frac{3c}{4}-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.29)
    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.30)
    and
    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{1+\left\|Ph_{1}\right\|^{2}\right\}$
  2. $2.$
    If $F$ is an anti-invariant Riemannian submersion (i.e. $J(\ker F_{\ast})\subseteq(\ker F_{\ast})^{\perp}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.32)
    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.33)
    and
    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{n+\left\|Ph_{1}\right\|^{2}\right\}$
  3. $3.$
    If $F$ is a semi-invariant Riemannian submersion (i.e. $J(\mathcal{D}_{1})=\mathcal{D}_{1}$ and $J(\mathcal{D}_{2})\subseteq(\ker F_{\ast})^{\perp}$), and
    1. $(i)$
[…]
@@ 2. $(ii)$
[…]
  4. $4.$
    If $F$ is a slant Riemannian submersion with angle $\theta$, then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)+\frac{3c}{4}\cos^{2}\theta-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.35)
    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.36)
    and
    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{n\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\cos^{2}\theta\right\}$
  5. $5.$
    If $F$ is a semi-slant Riemannian submersion (i.e. $J({\cal D}_{1})=\mathcal{D}_{1}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
  6. $6.$
    If $F$ is a hemi-slant Riemannian submersion (i.e. $J({\cal D}_{1})\subseteq(\ker F_{\ast})^{\perp}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
Remark 3.7.
In the above theorem, we observe that the corresponding inequalities become the same as in [23, Theorems 4 and 5], [22, Theorem 4.1], and [1, Equations (37), (38), and (62)]. Thus, the inequalities obtained using general forms are validated.
For generalized Sasakian space forms
Here, we obtain the Chen-Ricci inequalities for Riemannian submersions whose source spaces are generalized Sasakian space forms $\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)$ using the general inequalities obtained in Subsection 3.2. We also provide these inequalities when Riemannian submersions are geometrically structured and validate them with the existing particular cases where the source spaces are Sasakian, Kenmotsu, cosymplectic, and generalized Sasakian space forms. In this case, for any ${\cal Y}\in\Gamma(TM_{1})$, we write
$\phi{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ such that $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
Theorem 3.8.
Let $F:\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized Sasakian space form onto a Riemannian manifold. Then we have
@@
[…]
Proof.
Let $M_{1}$ be a generalized Sasakian space form. Then, using (2), we derive
@@
Then by using (3.49) and (3.11), (3.52) and (3.15), (3.3) and (3.3), we obtain the desired inequalities (3.8), (3.8) and (3.8) respectively. ∎
Remark 3.9.
[…]
From [28, 3, 18, 4, 56, 27], we have the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost contact metric manifolds onto Riemannian manifolds. These notions impose different-different conditions on the distribution $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the contact structure $\phi$. Using these conditions directly, we can prove the following result along the lines of the proof of Theorem 3.8.
Theorem 3.10.
Let $F:\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized Sasakian space form onto a Riemannian manifold.
  1. $1.$
    If $F$ is an invariant Riemannian submersion (i.e. $\phi(\ker F_{\ast})=\ker F_{\ast}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}$
    $\geq\left\{\begin{array}[]{l}3c_{2}\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 3c_{2},~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.61)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
    $-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}-c_{1}\left\{nr+n+r-2\right\}$
    $\geq\left\{\begin{array}[]{l}3c_{2}\left\{\left\|Ph_{1}\right\|^{2}+\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\right\}-c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\left\{1+\left\|Ph_{1}\right\|^{2}\right\}-c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.67)
  2. $2.$
    If $F$ is an anti-invariant Riemannian submersion (i.e. $\phi(\ker F_{\ast})\subseteq(\ker F_{\ast})^{\perp}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})$- $c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}$
    and
    $+\frac{n^{2}}{4}\left\|H\right\|^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
    $-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}$
    $-c_{1}\left\{nr+n+r-2\right\}-3c_{2}\left\{n-1+\left\|Ph_{1}\right\|^{2}\right\}$
  3. $3.$
    If $F$ is a semi-invariant Riemannian submersion (i.e. $\phi(\mathcal{D}_{1})=\mathcal{D}_{1}$ and $\phi(\mathcal{D}_{2})\subseteq(\ker F_{\ast})^{\perp}$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
  4. $4.$
    If $F$ is a slant Riemannian submersion with angle $\theta$, then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}\geq$
    $\left\{\begin{array}[]{l}3c_{2}\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\cos^{2}\theta-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\cos^{2}\theta,~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.79)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}$
    $+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}-c_{1}\left\{nr+n+r-2\right\}$
    $\geq\left\{\begin{array}[]{ll}3c_{2}\left\{\left(n-1\right)\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\cos^{2}\theta\right\}\\ -c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},&~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\left\{\left(n-1\right)\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\cos^{2}\theta\right\}\\ -c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},&~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.87)
  5. $5.$
    If $F$ is a semi-slant Riemannian submersion (i.e. $\phi({\cal D}_{1})=\mathcal{D}_{1}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
  6. $6.$
    If $F$ is a hemi-slant Riemannian submersion (i.e. $\phi({\cal D}_{1})\subseteq(\ker F_{\ast})^{\perp}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
Remark 3.11.
If we put the appropriate values of $c_{1}$ and $c_{2}=c_{3}$ from Table 4, then we observe that the corresponding inequalities reduce to the inequalities obtained in [9, 10, Theorems 4.1, 4.2, 4.3, 4.4, 4.5, and 4.6], [48, Theorems 4.1, 4.2, 4.3, 4.6, 4.7, and 4.8], [2, Equations (44), (53), (66), (73), and (74)], [49, Theorems 3.4, 3.5, and 3.6], and [7, Theorems 6, 7, 8, 9, and 10]. Thus, the inequalities obtained using general forms are validated.
@@ 4.1 Background
In this subsection, we collect some basic information that is needed for the remaining subsections.
@@
$\left(\nabla F_{\ast}\right)\left({\cal Z}_{1},{\cal Z}_{2}\right)=\nabla_{{\cal Z}_{1}}^{F}F_{\ast}({\cal Z}_{2})-F_{\ast}\left(\nabla_{{\cal Z}_{1}}^{{{1}}}{\cal Z}_{2}\right)$
for ${\cal Z}_{1},{\cal Z}_{2}\in\Gamma(T{M_{1}})$, where $\nabla^{F}$ is the pull-back connection. In addition, by [53] $\left(\nabla F_{\ast}\right)\left(X,Y\right)$ is completely contained in $\left({\rm range~}F_{\ast}\right)^{\perp}~{\rm for\ all}\ X,Y\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$.
@@
The Gauss equation for $F$ is defined as [51, p. 189]
@@
Supposing $\{V_{r+1},\dots,V_{m_{2}}\}$ an orthonormal basis of $\left({\rm range~}F_{\ast}\right)^{\perp}$ we set,
@@
In the sequel, for any ${\cal Z}\in\Gamma(TM_{2})$, we write
$J{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.5)
and
$\phi{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.6)
where $P{\cal Z}\in\Gamma({\rm range~}F_{\ast})$, $Q{\cal Z}\in\Gamma({\rm range~}F_{\ast})^{\perp}$ and $P$ is the endomorphism on $({\rm range~}F_{\ast})$ such that $\|PF_{\ast}h_{1}\|^{2}=\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}=\sum\limits_{j=1}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}$.
4.2 General framework of Chen-Ricci inequalities
In this subsection, we present a general form of the Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following theorem.
@@ 2. $(ii)$
At a point $p\in{M_{1}}$, let $\left\{h_{1},\dots,h_{r}\right\}$, $\{F_{\ast}h_{1},\dots,F_{\ast}h_{r}\}$ and $\{V_{r+1},\dots,V_{m_{2}}\}$ be orthonormal bases for $(\ker F_{\ast})^{\perp}$, $({\rm range~}~F_{\ast})$, and $({\rm range~}~F_{\ast})^{\perp}$, respectively. Then by putting $Z_{1}=Z_{4}=h_{i}$ and $Z_{2}=Z_{3}=h_{j}$ in (4.2), we obtain

### v2

@@ 3.4 Applications
In this subsection, we obtain the Chen-Ricci inequalities for Riemannian submersions from generalized complex and generalized Sasakian space forms. Toward this, we use the general forms established in Theorem 3.1. Also, for any ${\cal Y}\in\Gamma(TM_{1})$, we decompose
$J{\cal Y}{\textrm{~}or~}\phi{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ with $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
@@
  1. $(i)$
    If $(M_{1}(c_{1},c_{2}),g_{1},J)$ is a generalized complex space form, then
@@ 2. $(ii)$
    If $\left({M_{1}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{1}\right)$ is a generalized Sasakian space form, then
@@
In addition, the equality cases are similar to those of Theorem 3.1.
Proof of $(i)$.
Using (2.1), we derive
and
Adding equations (3.34), (3.35), and (3.36), we obtain
Then by using (3.34) and (3.11), (3.35) and (3.12), (3.4) and (3.1), we obtain the desired inequalities (3.22), (3.23) and (i) respectively. ∎
Proof of $(ii)$.
Using (2), we derive
@@
Then by using (3.40) and (3.11), (3.43) and (3.12), (3.4) and (3.1), we obtain the desired inequalities (ii), (ii) and (ii) respectively. ∎
@@ 4.1 Background
$\left(\nabla F_{\ast}\right)\left({\cal Z}_{1},{\cal Z}_{2}\right)=\nabla_{F_{\ast}{\cal Z}_{1}}^{2}F_{\ast}({\cal Z}_{2})-F_{\ast}\left(\nabla_{{\cal Z}_{1}}^{{{1}}}{\cal Z}_{2}\right)$ (4.2)
for ${\cal Z}_{1},{\cal Z}_{2}\in\Gamma(T{M_{1}})$. In addition, by [34] $\left(\nabla F_{\ast}\right)\left(X,Y\right)$ is completely contained in $\left({\rm range~}F_{\ast}\right)^{\perp}$ for all $X,Y\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$.
@@
The Gauss equation for $F$ is defined as [32, p. 189]
@@
Supposing $\{V_{r+1},\dots,V_{m_{2}}\}$ an orthonormal basis of $\left({\rm range~}F_{\ast}\right)^{\perp}$ we fix,
@@
4.2 General Chen-Ricci inequality
Here, we present general form of the Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following result.
@@ 2. $(b)$
Substituting $Z_{1}=Z_{4}=h_{i}$ and $Z_{2}=Z_{3}=h_{j}$ in (4.3), we obtain
@@
4.3 Improved general Chen-Ricci inequality
In this subsection, we present general form of the improved Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following result.
Theorem 4.2.
  1. $(i)$
    $B^{{\cal H}}=0$, equivalently, $F$ is totally geodesic, or
  2. $(ii)$
    $r=2$ and ( F_) ( h_1,h_1) = 3μV_r+1, ( F_) ( h_2,h_2) = μV_r+1, ( F_) ( h_1,h_2) =μV_r+2 for some function $\mu$ with respect to some orthonormal local frame field $\left\{h_{1},h_{2}\right\}$.
Proof.
For chosen bases, let
By (4.3), we get
Using (4.14) in (4.15), we derive
[…]
4.4 Example
Here, we give an example of a Riemannian map that satisfies the equality of Chen-Ricci and improved Chen-Ricci inequalities derived in Theorems 4.1 and 4.2, as follows.
Example 4.3.
Let
$\left(M_{1}=\left\{(x_{1},x_{2},x_{3},x_{4})\in\mathbb{R}^{4}:x_{i}\neq 0~\forall~1\leq i\leq 4\right\},g_{1}=e^{2x_{1}}\sum\limits_{i=1}^{2}dx_{i}^{2}+\sum\limits_{i=3}^{4}dx_{i}^{2}\right)$
and $\left(M_{2}=\left\{(y_{1},y_{2},y_{3},y_{4})\in\mathbb{R}^{4}\right\},g_{2}=e^{2y_{1}}\sum\limits_{j=1}^{2}dx_{j}^{2}+\sum\limits_{j=3}^{4}dx_{j}^{2}\right)$ be two Riemannian manifolds. Then by [29], $F:(M_{1},g_{1})\to(M_{2},g_{2})$ defined by
$F(x_{1},x_{2},x_{3},x_{4})=(x_{1},x_{2},x_{3},0)$
is a Riemannian map with
$(\ker F_{\ast})^{\perp}=\operatorname{span}\left\{h_{i}=\frac{\partial}{\partial x_{i}}\right\}_{i=1}^{3},\quad(\ker F_{\ast})=\operatorname{span}\left\{h_{4}=\frac{\partial}{\partial x_{4}}\right\},$
$({\rm range}~F_{\ast})=\operatorname{span}\left\{F_{\ast}h_{j}=\frac{\partial}{\partial y_{j}}\right\}_{j=1}^{3},\quad({\rm range}~F_{\ast})^{\perp}=\operatorname{span}\left\{\frac{\partial}{\partial y_{4}}\right\},$
[…]